Eliot needs to replace a window that is in the shape of a semi-circle.
What is the area of the window? Use 3.14
Round your answerto the nearest square inch.
radius equals 15in.

Answer by rfer(16322)

a=(pir^2)/2
a=(3.14*15^2)/2
a=706.5/2
a=353 sq in